# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v3

@@ 3 Partial recovery in correlated random graphs
Assuming Conjecture 2.5, there exists constant $C$ such that for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$, it holds that
@@
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ such that
@@
Without losing of generality, we may assume that $0\leq g_{i}\leq 1$, since otherwise we may replace $g_{i}$ with $\min\{\max\{g_{i},0\},1\}$, which will only make the left hand side of (3.6) smaller. Denote
@@
¯$\overline{\mathbb{P}}(g_{i}>\tfrac{c}{2}\mid\pi_{*}(1)=i)\geq\Omega(1)\,.$ (3.9)
@@
¯$\overline{\mathbb{P}}(\cdot\mid\pi_{*}(1)=i)\lhd_{\leq D}\mathbb{Q}\,.$
Thus we must have $\mathbb{Q}(g_{i}>\tfrac{c}{2})\geq\Omega(1)$. This yields that
${}\mathbb{Q}^{\prime}(g_{i}>\tfrac{c}{2})\geq\Omega(1)\mbox{ for all }\operatorname{TV}(\mathbb{Q}^{\prime},\mathbb{Q})=o(1)\,.$ (3.10)
This further suggests that
${}\mathbb{E}_{\mathbb{Q}^{\prime}}[g_{i}]=\Omega(1)\mbox{ for all }\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)\mbox{ and }i\in\Lambda\,.$ (3.11)
To this end, define $g=g_{1}+\ldots+g_{n}$. Using (3.8) and (3.11), we see that (recall again that $0\leq g_{i}\leq 1$)
${}\mathbb{E}_{\mathbb{Q}^{\prime}}[g]=\Omega(n)\mbox{ for all }\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)\,.$ (3.12)
However, using $0\leq g\leq n$, we see that for all $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ we have
$\mathbb{E}_{\mathbb{P}^{\prime}}[g^{2}]$ $\leq\operatorname{TV}(\overline{\mathbb{P}},\mathbb{P}^{\prime})\cdot\|g\|_{\infty}^{2}+\mathbb{E}_{\overline{\mathbb{P}}}[g^{2}]\leq o(n^{2})+2\big{(}1+\mathbb{E}_{\overline{\mathbb{P}}}[(g-1)^{2}]\big{)}$
$\leq o(n^{2})+2\mathbb{E}_{\overline{\mathbb{P}}}\Big{[}\Big{(}\sum_{i=1}^{n}(\mathbf{1}_{\{\pi_{*}(1)=i\}}-g_{i})\Big{)}^{2}\Big{]}$
@@
Combining (3.12) and (3.13), we see that for all $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime}),\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ we have
${}\frac{\mathbb{E}_{\mathbb{Q}^{\prime}}[g]}{\sqrt{\mathbb{E}_{\mathbb{P}^{\prime}}[g^{2}]}}=\omega(1)\,.$ (3.14)
Thus, using Item (2) in Conjecture 2.5 on the testing problem between $\mathbb{P}$ and $\mathbb{Q}$, we see that there exists an algorithm with running time $n^{D/(\log n)^{C}}$ that strongly distinguish $\mathbb{P}$ and $\mathbb{Q}$. However, it is impossible assuming Item (1) of Conjecture 2.5 as [DDL23+] showed that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$ where $\operatorname{TV}(\overline{\mathbb{P}},\mathbb{P})=o(1)$. This implies that (3.6) is impossible and thus completes our proof. ∎
Now we can finish the proof of Item (1) in Proposition 3.3.
Proof of Item (1) in Proposition 3.3.
We choose $C$ as in the statement of Lemma 3.6. Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. Using Items (1) and (2) in Definition 3.1, we see that $f_{j}f_{k}=0$ for all $k\neq j$, and thus
$1=\mathbb{E}_{\mathbb{P}}\big{[}(f_{1}+\ldots+f_{n})^{2}\big{]}=\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}}\big{[}f_{j}^{2}\big{]}\,.$ (3.15)
@@
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{\pi_{*}(i)}\big{]}=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{j}\mid\pi_{*}(i)=j\big{]}\geq c\,.$ (3.16)
@@
$\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{j}\big{)}^{2}\Big{]}=1-\Omega(1)\mbox{ where }g_{j}=\tfrac{1-\lambda}{n}+\lambda f_{j}\,,$
@@
This subsection is devoted to the proof of Item (2) in Proposition 3.3. Recall Definitions 1.3 and 1.4. Throughout this subsection, we will denote $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},\sigma_{*},G,A,B)$ where $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and $\mathbb{P}$ the marginal law of $(A,B)$. In addition, we denote $\mathbb{Q}$ to be the law of a pair of independent Erdős-Rényi models $\mathcal{G}(n,\tfrac{\lambda s}{n})$. In addition, we assume throughout this subsection that there exists a small constant $0.010<\delta<$ 0.01 such that
@@
$(\sqrt{\alpha}-\delta)(1+\epsilon^{N}k)\leq\sqrt{\alpha}-\delta/2\,;\quad 10k(1-\delta)^{N}\leq(1-\delta/2)^{N}\,;$ (3.18)
@@
where $J^{\prime}$ and $K^{\prime}$ are independent Bernoulli variables with parameter $s$. Let ~=~$\widetilde{\mathbb{P}}_{*}=\widetilde{\mathbb{P}}_{*,n}$ be the law of $(\sigma_{*},\pi_{*},G,G^{\prime},A^{\prime},B^{\prime})$ and denote ~=~$\widetilde{\mathbb{P}}=\widetilde{\mathbb{P}}_{n}$ the marginal law of $(A^{\prime},B^{\prime})$.
@@ 4 Detection in correlated SBMs
$\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta,k}(1)\mbox{ and }\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=O_{\delta,k}(1)\,.$
@@
=\ = $k^{-\frac{n}{2}}\mathbb{E}_{\mathbb{P}_{\sigma}}\Bigg{[}\prod_{(i,j)\in E(S)}\frac{G_{i,j}-\frac{\lambda}{n}}{\sqrt{\frac{\lambda}{n}(1-\frac{\lambda}{n})}}\cdot\prod_{(i,j)\in E(H)}\frac{G_{i,j}-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}}{\sqrt{\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}(1-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n})}}\Bigg{]}\,.$ (4.15)
Clearly, when $H\not\subset S$ we have (4.15) cancels to 00, since (denote $(i,j)$ to be an element in $E(H)\setminus E(S)$) under $\mathbb{P}_{\sigma}$ we have $\mathbb{E}_{\mathbb{P}_{\sigma}}[G_{i,j}-\tfrac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}]=0$ and all the edges $\{G_{i,j}:(i,j)\in\operatorname{U}_{n}\}$ are independent. In addition, from direct calculation we have that
@@
=\ = $\big{(}\tfrac{\epsilon^{2}\lambda}{n}\big{)}^{\frac{|E(S)|-|E(H)|}{2}}\sum_{\sigma_{\mathtt{V}}\in[n]^{\mathtt{V}}}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\times\sum_{\sigma_{[n]\setminus\mathtt{V}}\in[k]^{[n]\setminus\mathtt{V}}}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})=0\,.$
@@
Ξ()=1 and ${}\Xi(\emptyset)=1\mbox{ and }\Xi(S)=0\mbox{ for }\mathcal{L}(S)\neq\emptyset$ (4.21)
@@
×$\times\ \mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\Big{]}\,,$
@@
Acknowledgment. The author thanks Jian Ding and Jingqiu Ding for helpful discussions, Alexander S. Wein for helpful comments on the revised low-degree conjecture, and Hang Du for pointing him to the reference [SW25]. This work is partially supported by National Key R&\&&D program of China (No. 2023YFA1010103) and NSFC Key Program (Project No. 12231002).
@@ Appendix A Preliminary results in graphs and probability
In this section we include some preliminary results we will use in the main part of the paper. We first state some results on the structural properties of graphs established in [DDL23+, CDGL24+].
Lemma A.1 ([DDL23+], Lemma A.1).
@@
    #$\#\big{\{}T^{\prime}\Subset\mathcal{K}_{n}:S\Subset T^{\prime},|V(T^{\prime})|-|V(S)|=k,|E(T^{\prime})|-|E(S)|=l\big{\}}\leq n^{k}(|V(S)|+k)^{2l}$.
@@
    #$\#\big{\{}T^{\prime}\Subset S:|E(S)|-|E(T^{\prime})|=k\big{\}}\leq|E(S)|^{k}$.
@@
#$\#\Big{\{}S\mbox{ admissible}\!:H\ltimes S;|\mathfrak{C}_{l}(S,H)|=0\mbox{ for }l>N;|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H)=m;$
@@
Denote $v\in\mathcal{L}(S)\setminus V(H)$ and $(v,w)\in E(S)$. Define $\sigma_{U}$ and $\sigma_{\setminus U}$ to be the restriction of $\sigma$ on $U$ and on $[n]\setminus U$, respectively. Also define $\nu_{U}$ and $\nu_{\setminus U}$ to be the restriction of $\nu$ on $U$ and on $[n]\setminus U$, respectively. Then we have (let $\mathtt{V}=V(H)$)
@@
=\ = $\mathbb{E}_{\sigma_{([n]\setminus\mathtt{V})}\sim\nu_{([n]\setminus\mathtt{V})}}\Big{[}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\Big{]}$
=\ = $\mathbb{E}_{\sigma_{([n]\setminus(\mathtt{V}\cup\{v\}))}\sim\nu_{([n]\setminus(\mathtt{V}\cup\{v\}))}}\mathbb{E}_{\sigma_{\{v\}}\sim\nu_{\{v\}}}\Big{[}\omega(\sigma_{v},\sigma_{w})\prod_{(i,j)\in E(S)\setminus(E(H)\cup\{v,w\})}\omega(\sigma_{i},\sigma_{j})\Big{]}=0\,.\qed$
Appendix B Supplementary proofs
B.1 Proof of Theorem 2.4
The goal of this section is devoted to the proof of Theorem 2.4. We first briefly explain our proof ideas. Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ with
[…]
${}\mathbb{P}^{\prime}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)\mbox{ and }\mathbb{Q}^{\prime}(\mathcal{A}(\mathsf{Y})=0)=1-\epsilon\mbox{ for some }\epsilon=\epsilon_{n}\to 0\,.$ (B.2)
Define
${}M=M_{n}=\max\big{\{}n,\lceil\epsilon_{n}^{-\frac{1}{2}}\rceil\big{\}}\to\infty\,.$ (B.3)
Definition B.1.
  • ¯$\overline{\mathcal{H}}_{0}$: we let $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M}$ to be independently sampled from $\mathbb{Q}^{\prime}$.
  • ¯$\overline{\mathcal{H}}_{1}$: we first sample $\kappa\in\{1,\ldots,M\}$ uniformly at random, and (conditioned on the value of $\kappa$) we let $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M}$ are independent samples with $\mathsf{Y}_{\kappa}$ generated from $\mathbb{P}^{\prime}$ and $\{\mathsf{Y}_{j}:j\neq\kappa\}$ generated from $\mathbb{Q}^{\prime}$.
In addition, denote $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$ to be the law of $(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})$ under ¯$\overline{\mathcal{H}}_{1}$ and ¯$\overline{\mathcal{H}}_{0}$, respectively.
Now assuming that (B.2) holds, we see that
[…]
¯$\overline{\mathbb{P}}\Big{(}\big{(}\mathcal{A}(\mathsf{Y}_{1}),\ldots,\mathcal{A}(\mathsf{Y}_{M})\big{)}\neq(0,\ldots,0)\Big{)}\geq\Omega(1)\,.$ (B.5)
Lemma B.2.
Proof.
Note that
$\frac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\overline{\mathbb{P}}(\cdot\mid\kappa=i)}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}(\mathsf{Y}_{i})\,.$ (B.6)
In addition, recall (2.1). Define
$\mathcal{N}_{D}=\big{\{}f\in\mathcal{P}_{D}:f=0\mbox{ a.s. under }\mathbb{Q}^{\prime}\big{\}}\,.$
${}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=\Bigg{(}\sum_{\alpha\in\Lambda}\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}(\mathsf{Y})]^{2}\Bigg{)}^{\frac{1}{2}}\,.$ (B.7)
For any $f\in\mathcal{N}_{D}^{\perp}$, it can be uniquely expressed as
$f=\sum_{\alpha\in\Lambda}C_{\alpha}f_{\alpha}\,,$
with equality holds if and only if $C_{\alpha}\propto\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}]$. This yields (B.7). Now, note that ¯$\overline{\mathbb{Q}}=(\mathbb{Q}^{\prime})^{\otimes M}$ is a product measure of $\mathbb{Q}^{\prime}$, there is a natural standard orthogonal basis under $\overline{\mathbb{Q}}$, given by
Thus, similarly as in (B.7), we see that
$\Big{(}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}\big{)}\Big{)}^{2}=\sum_{\begin{subarray}{c}(\alpha_{1},\ldots,\alpha_{M}):\alpha_{i}\in\Lambda\\ \sum_{i=1}^{M}\operatorname{deg}(f_{\alpha_{i}})\leq D\end{subarray}}\mathbb{E}_{\overline{\mathbb{P}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\Big{]}^{2}\,.$ (B.8)
In addition, using (B.6), we see from direct calculation that
=(B.6)${=}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}_{\overline{\mathbb{Q}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\cdot\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}(\mathsf{Y}_{i})\Big{]}$
=$=\begin{cases}1\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0)\,;\\ \frac{1}{M}\mathbb{E}_{\mathbb{P}^{\prime}}\big{[}f_{\alpha_{j}}(\mathsf{Y}_{j})\big{]}\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0,\alpha_{j},0,\ldots,0)\,;\\ 0\,,&\mbox{otherwise}\,.\end{cases}$ (B.9)
Plugging (B.9) into (B.8), we get that
@@ Appendix B Supplementary proofs
where in the second equality we use (B.7) and the assumption that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$. This completes our proof. ∎
We can now finish the proof of Theorem 2.4.
Proof of Theorem 2.4.
Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ satisfying (B.1). Consider the hypothesis testing problem stated in Definition B.1. Using (B.4) and (B.5), we see that there is an efficient algorithm that achieves weak detection between $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$, which contradicts with Lemma B.2 and the equivalence between Items (1) and (2) in Conjecture 2.2. ∎
B.2 Proof of Lemma 3.5

### v4

@@ 3 Partial recovery in correlated random graphs
Assuming Conjecture 2.2, there exists a constant $C$ such that for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ such that each $g_{i,j}(A,B)$ can be computed in time $n^{D/(\log n)^{C}}$ and
[…]
it holds that
@@
Without loss of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ such that $\sum_{1\leq j\leq n}g_{j}(A,B)=o(n)$ for all graphs $A,B$ and
@@
Without loss of generality, we may assume that $0\leq g_{i}\leq 1$, since otherwise we may replace $g_{i}$ with $\min\{\max\{g_{i},0\},1\}$, which will only make the left hand side of (3.10) smaller. Denote
@@
$\overline{\mathbb{P}}(g_{i}>\tfrac{c}{2}\mid\pi_{*}(1)=i)\geq\Omega(1)\,.$ (3.13)
@@
$\overline{\mathbb{P}}(\cdot\mid\pi_{*}(1)=i)\lhd_{\leq D}\mathbb{Q}\,.$
Thus we must have
$\mathbb{Q}(g_{i}>\tfrac{c}{2})\geq\Omega(1)\,.$ (3.14)
To this end, define $g=g_{1}+\ldots+g_{n}$. Using $0\leq g\leq n$, we see that (3.14) yields that $\mathbb{E}_{\mathbb{Q}}[g]=\Omega(n)$, which violates our assumption (3.8). This implies that (3.10) is impossible and thus completes our proof. ∎
Now we can finish the proof of Item (1) in Proposition 3.4 assuming Conjecture 2.2.
Proof of Item (1) in Proposition 3.4 assuming Conjecture 2.2.
Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ satisfying Items (1) and (2) in Definition 3.1 with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0<c<0.01$. We first claim that we may assume that
@@
In fact, denote
we may consider the truncated statistics
$f_{i,j}^{\prime}(A,B)=f_{i,j}(A,B)\cdot\mathbf{1}_{\mathcal{A}}\,.$
It is straightforward to check that $\{f_{i,j}^{\prime}(A,B):1\leq i,j\leq n\}$ satisfies (3.15) and we still have
$\mathbb{E}_{\mathbb{P}_{*}}\big[f_{i,j}^{\prime}(A,B)\big]\geq 1-c\,.$
Thus, we may assume that (3.15) holds without loss of generality. To this end, using (3.15), we see that $f_{1,j}f_{1,k}=0$ for all $k\neq j$, and thus
$1=\mathbb{E}_{\mathbb{P}}\big[(f_{1,1}+\ldots+f_{1,n})^{2}\big]=\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}}\big[f_{1,j}^{2}\big]\,.$ (3.16)
@@
$\mathbb{E}_{\mathbb{P}_{*}}\big[f_{1,\pi_{*}(i)}\big]=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\big[f_{1,j}\mid\pi_{*}(i)=j\big]\geq c\,.$ (3.17)
@@
$\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big[\big(\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{j}\big)^{2}\Big]=1-\Omega(1)\mbox{ where }g_{j}=\tfrac{1-\lambda}{n}+\lambda f_{1,j}\,$,
@@
This subsection is devoted to the proof of Item (2) in Proposition 3.4. Recall Definitions 1.3 and 1.4. Throughout this subsection, we will denote $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},\sigma_{*},G,A,B)$ where $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and $\mathbb{P}$ the marginal law of $(A,B)$. In addition, we denote $\mathbb{Q}$ to be the law of a pair of independent Erdős-Rényi models $\mathcal{G}(n,\tfrac{\lambda s}{n})$. In addition, we assume throughout this subsection that there exists a small constant $0<\delta<0.01$ such that
@@
[…]
@@
where $J^{\prime}$ and $K^{\prime}$ are independent Bernoulli variables with parameter $s$. Let $\widetilde{\mathbb{P}}_{*}=\widetilde{\mathbb{P}}_{*,n}$ be the law of $(\sigma_{*},\pi_{*},G,G^{\prime},A^{\prime},B^{\prime})$ and denote $\widetilde{\mathbb{P}}=\widetilde{\mathbb{P}}_{n}$ the marginal law of $(A^{\prime},B^{\prime})$.
@@ 4 Detection in correlated SBMs
[…]
@@
[…]
Clearly, when $H\not\subset S$ we have (4.15) cancels to 0, since (denote $(i,j)$ to be an element in $E(H)\setminus E(S)$) under $\mathbb{P}_{\sigma}$ we have $\mathbb{E}_{\mathbb{P}_{\sigma}}[G_{i,j}-\tfrac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}]=0$ and all the edges $\{G_{i,j}:(i,j)\in\operatorname{U}_{n}\}$ are independent. In addition, from direct calculation we have that
@@
$=\$$\big(\tfrac{\epsilon^{2}\lambda}{n}\big)^{\frac{|E(S)|-|E(H)|}{2}}\sum_{\sigma_{\mathtt{V}}\in[n]^{\mathtt{V}}}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\times\sum_{\sigma_{[n]\setminus\mathtt{V}}\in[k]^{[n]\setminus\mathtt{V}}}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})=0\,.$
@@
$\Xi(\emptyset)=1\mbox{ and }\Xi(S)=0\mbox{ for }\mathcal{L}(S)\neq\emptyset$ (4.21)
@@
$\times\ \mathbb{E}_{\sigma\sim\nu}\Big[\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\Big]\,$,
@@
Acknowledgment. The author thanks Jian Ding and Jingqiu Ding for helpful discussions, Alexander S. Wein for helpful comments on the revised low-degree conjecture, and Hang Du for pointing him to the reference [SW25].
@@ Appendix A Preliminary results in graphs and probability
In this section we include some preliminary results we will use in the main part of the paper. We first state some results on the structural properties of graphs established in [DDL25, CDGL24+].
Lemma A.1 ([DDL25], Lemma A.1).
@@
    $\big\{T^{\prime}\Subset\mathcal{K}_{n}:S\Subset T^{\prime},|V(T^{\prime})|-|V(S)|=k,|E(T^{\prime})|-|E(S)|=l\big\}\leq n^{k}(|V(S)|+k)^{2l}$.
@@
    $\big\{T^{\prime}\Subset S:|E(S)|-|E(T^{\prime})|=k\big\}\leq|E(S)|^{k}$.
@@
$\#\Big\{S\mbox{ admissible}\!:H\ltimes S;|\mathfrak{C}_{l}(S,H)|=0\mbox{ for }l>N;|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H)=m;$
@@
Denote $v\in\mathcal{L}(S)\setminus V(H)$ and $(v,w)\in E(S)$. Define $\sigma_{U}$ and $\sigma_{\setminus U}$ to be the restriction of $\sigma$ on $U$ and on $n]\setminus U$, respectively. Also define $\nu_{U}$ and $\nu_{\setminus U}$ to be the restriction of $\nu$ on $U$ and on $n]\setminus U$, respectively. Then we have (let $\mathtt{V}=V(H)$)
@@
Appendix B Supplementary proofs
B.1 Proof of Lemma 3.6